Find the equation of a line containing the given point and having the given slope. (10,8), m=13/20
Answer by VictoryStar(6)

You did not specify which form to use, therefore, I will explain all three.
To begin, we will use pointslope form, as the other forms will stem from this. Point slope is used for when you do not know the Y-intercept, or you are able to get the slope and a point(s). The form is written in y – y1 = m(x – x1). The y1 and x1 are only to place your points in, nothing else, where m, is simply your slope.
Your 8 is your y, (x,y), therefore, you start with y-8 = m(x - x1)
13/20, is your m, also known as slope, giving you y - 8 = 13/20 (x - x1)
10, is your x (x.y), making it so that your equation is y - 8 = 13/20 (x - 10)


Next is slope intercept form, where it is written in the form of Y=Mx+b. Since you have most of the needed information, simplifying the prior equation may give you the answer needed.
Start with distributing. (This is where you multiply a number outside of a parenthesis into all the inside numbers.) You will obtain y - 8 = 13x/20 - 13/2.
Add 8 to both sides, to get Y on it's own. (Also known as isolation.)
You will end up with Y= 13x/20 + 3/2 (-13/2 + 8 = -13/2 + 16/2 = 16/2 - 13/2 = 3/2)


Finally is standard form. There is no practical use of using it here, but I will go through it nevertheless. Simply get both Y and X on the left side. By subtracting 13x/20 to both sides, you will obtain
-13x/20 + Y = 3/2. But however, standard form requires that both numbers preceding Y and X cannot be fractions, therefore, you will multiply all by 20.
You will obtain -13x + 20 y = 15.
Some teachers do not approve of having the before number of x being negative, therefore, you may multiply each number by -1. (Simply reversing the signs.)
Following this process you will have the equation of 13x - 20y = -15.